lwaysRain Irrigation, Inc. would like to determine capacity requirements for the next four years. Currently, two production lines are in place for making bronze and plastic sprinklers. Three types of sprinklers are available in both bronze and plastic: 90-degree nozzle sprinklers, 180-degree nozzle sprinklers, and 360-degree nozzle sprinklers. Management has forecast demand for the next four years as follows:

YEARLY DEMAND
1 (IN 000s) 2 (IN 000s) 3 (IN 000s) 4 (IN 000s)
Plastic 90 32 46 54 58
Plastic 180 14 15 13 16
Plastic 360 47 54 63 67
Bronze 90 9 8 8 19
Bronze 180 4 4 9 15
Bronze 360 8 12 17 22


Both production lines can produce all the different types of nozzles. The bronze machines needed for the bronze sprinklers require two operators and can produce up to 20,000 sprinklers. The plastic injection molding machine needed for the plastic sprinklers requires three operators and can produce up to 320,000 sprinklers. Three bronze machines and only one injection molding machine are available.

What are the capacity requirements for the next four years? (Assume that there is no learning.) (Enter the demand values in thousands. Round your answers to 2 decimal places.)
Answer #1

Year I

Plastic machines needed = (32+14+47)*1000 /320000 = 0.2906 =1 machines

Bronze machies needed = (9+4+8)*1000 / 20000 =1.05 =2 machines

Yeas 2

Plastic machines needed = (46+15+54)*1000 /320000 = 0.3593 =1 machines

Bronze machies needed = (8+4+12)*1000 / 20000 = =1.2 =2 machines

Year 3

Plastic machines needed = (54+13+63)*1000 /320000 = 0.4062 =1 machines

Bronze machies needed = (8+9+17)*1000 / 20000 =1.7 =2 machines

Year 4

Plastic machines needed = (58+16+67)*1000 /320000 = 0.4406 =1 machines

Bronze machies needed = (19+15+22)*1000 / 20000 =2.8 =3 machines

It can be seen that no machines need to be purchased in next four years